OpenStudy (anonymous):

Can anyone help me understand solving systems by elimination? My online class does a uhhh, not-so-good job of explaining it... :)

OpenStudy (anonymous):

Do you mean systems of equations?

OpenStudy (anonymous):

yes! :/

OpenStudy (blacksteel):

The idea behind solving systems by elimination is to eliminate one variable in order to find the value of the other. For example, consider the system: x + y = 2 x - y = 4 If we add these two equations together, we get x + y + x - y = 2 + 4 2x = 6 x = 3 Then we can substitute this into one of the above equations to find y: 3 + y = 2 y = -1

OpenStudy (anonymous):

Eliminate a variable by making it's coefficient same in all equations you have, and then adding them, or subtracting one from the other. This should leave you with fewer variables. Do this until only one variable is left.

OpenStudy (anonymous):

it can be explained by an example solve the equations or x and y x + y = 10................(1) 2x -y = 8...................(20 we can add/subtract equations just as we can add/sub numbers and algebraic terms if we add (1) + (2) we eliminate y (y + (-y) = 00 so we get 3x = 18 x = 6 and we cn get value of x by plugging x = 6 in either (1) or (2)

OpenStudy (blacksteel):

Now, not all equations will have x and y coefficients that make elimination easy. The above problem worked well because the y terms had the same coefficients but with opposite signs, so when the two equations were added together they just canceled out. Suppose we have a system where this isn't the case, like: 3x + 7y = 1 x + 2y = 15 To solve this, we want to multiply both sides of one of the two equations by a number in order to make the coefficients of either the x's or y's the same. In this case, let's multiply the bottom equation by 3. Then we have: 3x + 7y = 1 3x + 6y = 45 Now we can subtract the equations from one another to get: 3x + 7y - 3x - 6y = 1 - 45 y = -44 Then we substitute this into our equation to get: x - 88 = 15 x = 103

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/Alg/SystemsTwoVrble.aspx Elimination method half way down page.

OpenStudy (anonymous):

Thanks EVERYONE! :D :D :D :D :D

OpenStudy (anonymous):

Help on my latest question please!??!?!? "Elimination!!! Solving systems of equations... Can anyone help me figure out these, using elimination? 3x-6y=3 7x-5y=-1 Thank yous! "

OpenStudy (blacksteel):

You should be able to solve this using the same technique given here. Either multiply one equation by a number that will give both equations the same coefficients for one variable (for example, multiply the first equation by 7/3) or multiply both equations by numbers that will do this (for example, the first equation by 7 and the second by 3)